write the equation in slope intercept form of the line that passes through (5,4) and (-6,1)
Answer by stanbon(75887)

write the equation in slope intercept form of the line that passes through (5,4) and (-6,1)
--------------------
slope = (4-1)/(5--6) = 3/11
-----------------------------
intercept ?
You know y=4 when x = 5, and m = 3/11 ; solve for "b"
4 = (3/11)*5 + b
4 = 15/11 + b
b = 44/11 - 15/11
b = 29/11
-----------------
Equation:
y = (3/11)x + (29/11)
=========================
Cheers,
Stan H.